A POINT IN SPACE AN OBJECT OF MASS 5KG EXPERIENCES AN ACCLERATON OF 4M/S2 HOW MUCH ACCLERATION WILL AN OBJECT OF MASS 1O KG UNDERGO?

Grade:9

1 Answers

BALA SRIVATSAV YERRAMILLI
3 years ago
Hi, here’s your answer.
Since F = ma, the force in the initial case will be F = 5 x 4 = 20N.
In the second case, the mass is now 10kg so, acceleration will be a = F/m = 20/10 = 2m/s^2
Hope this helps you!